a consider the normal distribution of IQs with a mean of 100 and a standard deviation of 16. What






percentage of IQs are?






a. greater than 95?






b. less than 120






c. between 90 and 110

1
Expert's answer
2022-03-18T16:36:12-0400

Population mean "\mu=100"

Standard deviation "\sigma=16"


(a) Greater than 95

"Z=\dfrac{X-\mu}{\sigma}"


"Z=\dfrac{95-100}{16}=-0.3125"


For values greater than 95

"P(1-Z)=1-0.37828"

"=0.62172"


Percentage

"=0.62172\times 100 =66.172\%"


(b) Less than 120

"Z=\dfrac{X-\mu}{\sigma}=\dfrac{120-100}{16}=1.25"


"P(Z=1.25)=0.89435"


Percentage

"=0.89435\times100=89.435\%"


(c) Between 90 and 110

"=\dfrac{90-100}{16}<Z<\dfrac{110-100}{16}"


"=P(-0.625<Z<0.625)"


"=P(Z<0.625)-P(Z<-0.625)"


"=0.73237-0.26763"


"=0.46474"


Percentage

"=0.46474\times100=46.474\%"